# Observational Auditing of Label Privacy

### v2

@@ 6 Experiments on auditing Label DP
We validate our theoretical framework with experiments on two representative datasets: CIFAR-10 (cifar10) for image classification and Criteo (criteo) for tabular data with sensitive labels. For each dataset, we train classifiers using standard Label DP mechanisms and empirically evaluate the privacy guarantees using our audit procedure. Appendix 12 shows similar results for a third large text dataset. We also evaluate our Label DP auditing technique for Randomized Response (warner1965randomized) on synthetic data.
@@ 6.1 Label DP learning algorithms
The earliest approach to achieving Label DP is the Randomized Response (RR) mechanism (warner1965randomized). In RR, each training label is randomly replaced according to a fixed probability distribution before being shared with the learning algorithm. This randomization helps protect the privacy of individual labels. We briefly review several recent Label DP mechanisms that improve on RR.
Label Private One-Stage Training (LP-1ST, google-labeldp) Instead of using a fixed distribution as in RR, LP-1ST samples each training label $y_{i}$ from a learned prior distribution $P(y\mid X_{i})$. The prior can be estimated by observing the top-$k$ predictions from a pretrained model (either in domain or out-of-domain), restricting RR to the most probable labels. Alternatively, the training can be split into multiple stages, where an earlier model provides the prior for the next (LP-MST).
Private Aggregation of Teacher Ensembles with FixMatch (PATE-FM, meta-labeldp) PATE-FM combines the FixMatch semi-supervised learning algorithm (sohn2020fixmatch) with private aggregation. Multiple teacher models are trained, each using all unlabeled examples and disjoint subsets of the labeled data. The predictions from these teachers are then aggregated in a differentially private manner using the PATE framework (papernot2016semi) to train a student model.
Additive Laplace with Iterative Bayesian Inference (ALIBI, meta-labeldp) ALIBI achieves Label DP by adding Laplace noise to the one-hot encoded labels (ghosh2012universally), making the released labels differentially private. The model is trained on soft labels computed from noisy observations via Bayesian inference.
@@ 6.4 Criteo experiments
The Criteo dataset contains user click-through data with demographic information encoded as 13 numerical features and 26 categorical features. A binary label indicates whether the user clicked on the ad. The distribution of the labels is highly imbalanced, with only 3% of positives (clicks). The overall dataset contains over 4 billion click log data points over a period of 24 days. We followed the same setup in wu2022does where 1 million data points are selected for each day. We divide the data into 80% for training, 4% for validation, and 16% for testing. Model performance is evaluated using the log-loss metric on the test set.
We train gradient boosting decision trees with the CatBoost library (dorogush2018catboost) with varying privacy budgets ${\epsilon}\in\{0.1,1,2,4,8,\infty\}$ and evaluate label privacy using observational privacy auditing (Table 2). Table 5 (Appendix 11) shows model performance under different Label DP algorithms and values $\epsilon$}.
@@
Table 2: Criteo. Auditing Label DP algorithms for different $\epsilon$} with $\delta=10^{-5}$. As in wu2022does, for LP-1ST (domain prior) at ${\epsilon}\in\{0.1,1,2\}$ the training process did not produce meaningful outcomes.
@@ 6.5 Comparison with Existing Methods
We compare our observational auditing approach against traditional canary-based methods to demonstrate the effectiveness and practicality of our framework. More specifically, we evaluate against the lightweight difficulty calibration MIA in watson2022difficulty, where membership scores are adjusted to the difficulty of correctly classifying the target sample. For each canary datapoint, we set the calibrated membership scores as the difference in the loss between the target model and the reference model $M^{\prime}$. Fig. 1 shows how our method achieves similar auditing results when compared to MIA on the CIFAR-10 and Criteo datasets.
@@
We leave to future work a comparison with more computationally intensive methods, such as zarifzadeh2024low, which require training multiple auxiliary (shadow) models to achieve state-of-the-art attack performances. However, we compare our approach to that of zarifzadeh2024low for the special case of one single reference model in Appendix 12.
Gap between theoretical and empirical epsilon. Fig. 1 shows a large gap between the theoretical and empirical bounds on $\epsilon$, especially for larger values of $\epsilon$. This gap is due to several factors that have been previously explored in the literature: (1) the attacks are black-box, i.e., they do not have access to intermediate outputs of the training procedure such as gradients (nasr2021adversary), (2) the theoretical $\epsilon$ is an upper bound that is not necessarily tight, (3) $f$-DP auditing is not always tight, especially at higher values of $\epsilon$} (mahloujifar2025auditing).

### v3

@@ 6 Experiments on auditing Label DP
We validate our theoretical framework with experiments on two representative datasets: CIFAR-10 (Krizhevsky, 2009) for image classification and Criteo (Criteo AI Lab, 2015) for tabular data with sensitive labels. For each dataset, we train classifiers using standard Label DP mechanisms and empirically evaluate the privacy guarantees using our audit procedure. Appendix 12 shows similar results for a third large text dataset. We also evaluate our Label DP auditing technique for Randomized Response (Warner, 1965) Our attack is implemented in the PrivacyGuard library.  [^1] https://github.com/facebookresearch/PrivacyGuard
@@ 6.1 Label DP learning algorithms
The earliest approach to achieving Label DP is the Randomized Response (RR) mechanism (Warner, 1965). In RR, each training label is randomly replaced according to a fixed probability distribution before being shared with the learning algorithm. This randomization helps protect the privacy of individual labels. We briefly review several recent Label DP mechanisms that improve on RR.
Label Private One-Stage Training (LP-1ST, Ghazi et al. (2021)) Instead of using a fixed distribution as in RR, LP-1ST samples each training label $y_{i}$ from a learned prior distribution $P(y\mid X_{i})$. The prior can be estimated by observing the top-$k$ predictions from a pretrained model (either in domain or out-of-domain), restricting RR to the most probable labels. Alternatively, the training can be split into multiple stages, where an earlier model provides the prior for the next (LP-MST).
Private Aggregation of Teacher Ensembles with FixMatch (PATE-FM, Malek et al. (2021)) PATE-FM combines the FixMatch semi-supervised learning algorithm (Sohn et al., 2020) with private aggregation. Multiple teacher models are trained, each using all unlabeled examples and disjoint subsets of the labeled data. The predictions from these teachers are then aggregated in a differentially private manner using the PATE framework (Papernot et al., 2017) to train a student model.
Additive Laplace with Iterative Bayesian Inference (ALIBI, Malek et al. (2021)) ALIBI achieves Label DP by adding Laplace noise to the one-hot encoded labels (Ghosh et al., 2012), making the released labels differentially private. The model is trained on soft labels computed from noisy observations via Bayesian inference.
@@ 6.4 Criteo experiments
The Criteo dataset contains user click-through data with demographic information encoded as 13 numerical features and 26 categorical features. A binary label indicates whether the user clicked on the ad. The distribution of the labels is highly imbalanced, with only 3% of positives (clicks). The overall dataset contains over 4 billion click log data points over a period of 24 days. We followed the same setup in Wu et al. (2023) where 1 million data points are selected for each day. We divide the data into 80% for training, 4% for validation, and 16% for testing. Model performance is evaluated using the log-loss metric on the test set.
We train gradient boosting decision trees with the CatBoost library (Prokhorenkova et al., 2018) with varying privacy budgets ${\epsilon}\in\{0.1,1,2,4,8,\infty\}$ and evaluate label privacy using observational privacy auditing (Table 2). Table 5 (Appendix 11) shows model performance under different Label DP algorithms and values $\epsilon$}.
@@
Table 2: Criteo. Auditing Label DP algorithms for different $\epsilon$} with $\delta=10^{-5}$. As in Wu et al. (2023), for LP-1ST (domain prior) at ${\epsilon}\in\{0.1,1,2\}$ the training process did not produce meaningful outcomes.
@@ 6.5 Comparison with Existing Methods
We compare our observational auditing approach against traditional canary-based methods to demonstrate the effectiveness and practicality of our framework. More specifically, we evaluate against the lightweight difficulty calibration MIA in Watson et al. (2022), where membership scores are adjusted to the difficulty of correctly classifying the target sample. For each canary datapoint, we set the calibrated membership scores as the difference in the loss between the target model and the reference model $M^{\prime}$. Fig. 1 shows how our method achieves similar auditing results when compared to MIA on the CIFAR-10 and Criteo datasets.
@@
We leave to future work a comparison with more computationally intensive methods, such as Zarifzadeh et al. (2024), which require training multiple auxiliary (shadow) models to achieve state-of-the-art attack performances. However, we compare our approach to that of Zarifzadeh et al. (2024) for the special case of one single reference model in Appendix 12.
Gap between theoretical and empirical epsilon. Fig. 1 shows a large gap between the theoretical and empirical bounds on $\epsilon$, especially for larger values of $\epsilon$. This gap is due to several factors that have been previously explored in the literature: (1) the attacks are black-box, i.e., they do not have access to intermediate outputs of the training procedure such as gradients (Nasr et al., 2021), (2) the theoretical $\epsilon$ is an upper bound that is not necessarily tight, (3) $f$-DP auditing is not always tight, especially at higher values of $\epsilon$} (Mahloujifar et al., 2025b).